Homework: HW 3 (Chapter 4) Score: 0 of 1 pt Problem 4.12 Hw Score: 16.67%, 1 of 6 pts Question Help * Consider the following actual and forecast demand levels for Big Mac Actual Demand 88.00 75.00 68.00 52.00 Forecast Demand 88.00 88.00 84.75 80.56 Day Tuesday Wednesday Thursday Friday The forecast for Monday was derived by obs forecasts were derived by using exponential smoothing with a smoothing constant of 0.25. Using this exponential smoothing method, the forecast for Big Mac demand for Friday isBig Macs (round your response to one decimal place) All parts showing t
media%2Faa0%2Faa0002ad-b078-4d42-8638-41
media%2F6b0%2F6b0752db-4f14-4ca0-b4ca-c2
Answer #1

Note: Please rate the answer.

Answer: 4.12

  • 73.42 (Answer)

Explanation:

  • In exponential smoothing:
  • Ft+1 = Ft + \alpha (At -Ft)
  • Where Ft+t is the forecast of cuurent period T+1, based on forcast and actual demand of previous periods (Ft and At respectively).
  • Ffriday = Fthursdays + \alpha ( Athursdays -Fthursdays) = 80.56 + 0.25 * (52 -80.56) = 73.42

Answer: 4.14

  • 0.085

Explanation:

  • MAD = \frac{1}{4}\sum_{i=1}^{4} \left | Actual - forecast \right |_{i}
  • Solving MAD for method 1 = (0.13+0+0.05+0.16)/ 4 =0.085

Answer 4.37 (a)

  • 2.5

Explanation:

  • MAD = \frac{1}{4}\sum_{i=1}^{4} \left | Actual - forecast \right |_{i}
  • Solving MAD for method 1 = (5+2+3+0)/ 4 =2.5